[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Fri Nov 16 08:55:02 EST 2018

Logentry Text:
RC report:
At ~17 hrs beam was restored to the Hall and we began taking data for kinematic group 9 with 60 uA on LH2.  However, after ~30 minutes of 55 uA, the 1L27 fast valve would not open (https://logbooks.jlab.org/entry/3626790) and experts had to be called in.  Beam was restored around 19:20 and data collection continued with kinematic group 9-3.  

Beam throughout swing was rather unreliable (https://logbooks.jlab.org/entry/3627075) as a result of mismatches in arc 9 (https://logbooks.jlab.org/entry/3626853), repeated valve closures, burping, and vacuum issues in 1L26 and 1L27 (https://logbooks.jlab.org/entry/3626823, https://logbooks.jlab.org/entry/3626855, https://logbooks.jlab.org/entry/3626881), injector mismatches (https://logbooks.jlab.org/entry/3626861), and a lost IOC communication (https://logbooks.jlab.org/entry/3626896).  At 00:21 a vacuum issue in 1L26 took beam down for a little more than 2 hrs.  At 00:20 hrs the SHMS HB coil 1 quenched and subsequently fast dumped (https://logbooks.jlab.org/entry/3626924, https://logbooks.jlab.org/entry/3626927).  The interlocks and quench detector were reset and power supply cycled once the current reached 0 A.  The HB current was then cycled to the MOL current and remained stable for the remainder of the evening.

Once beam was restored data collection continued throughout the remainder of owl and the beginning of day shift and 3 kinematic settings in group 9 were completed.  We will likely remain in kinematic group 9 until owl or day shift Saturday at which time we will transition into kinematic group 13.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Planned accesses to the hall:
Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K.

Special requests:
Vijay Kumar (Regina group) has requested a walkthrough during the next access
Jay B. would like to conduct another beam steering test (5 minute procedure) during the next polarity change.  He requests to only be contacted if the polarity change occurs during the hours of 0630 and 1930.
Carlos Munoz Camacho would like access to the hall to show Hall-C to two French engineers  working on the NPS project.  It should be a ~30 minute access, and is a low priority.  He can be contacted at 357-1065

We will meet today at 4:30 pm in the second floor of the counting house.

* Bluejeans remote connection information:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial:
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3627122
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181116/8ef730f0/attachment.html>

More information about the Hallc_running mailing list